(a) How much heat is required to raise by 12.40°C the temperature of 1.8 kg of water. (b) How much heat is required to raise by 12.40°C the temperature of 1.8 kg of iron.
Answer #1

(a) heat required = mc*dT

where m = mass

c= specific heat capacity = 4186 J/kg K

dT = change in temperature

Q = 1.8*4186*12.4

= 93431.52J

(b) specific heat capacity of iron = 450 J/kg K

Q = 1.2*450*12.4

= 10044 J